OpenStudy (anonymous):

What value of k solves the equation? k^-3=1/27 A. –81 B. –9 C. –3 D. 3 -3 and 3 have the same answers so I'm kind of stuck...

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (mathstudent55):

\(\large k^{-3} = \dfrac{1}{27} \) Do you understand negative exponents?

OpenStudy (mathstudent55):

For example, what is \(\large 5^{-2}\) ?

OpenStudy (anonymous):

0.04?

OpenStudy (mathstudent55):

Yes, that is correct, but my question is without using a calculator, how do you deal with a negative exponent?

OpenStudy (mathstudent55):

In other words, what is \(\large 5^{-2} \) equal to using a positive exponent.

OpenStudy (anonymous):

4%

OpenStudy (mathstudent55):

Here it is. The general rule is: \(\large a^{-n} = \dfrac{1}{a^n} \) Here's the example I asked about above: \( \large 5^{-2} = \dfrac{1}{5^2} \)

OpenStudy (mathstudent55):

When you raise a number to a negative exponent, it means you write 1 over the number raised to the same but positive exponent.

OpenStudy (mathstudent55):

Here are a few more examples: \(\large 3^{-5} = \dfrac{1}{3^5} \) \(\large x^{-6} = \dfrac{1}{x^6} \) \(\large y^{-a} = \dfrac{1}{y^a} \)

OpenStudy (mathstudent55):

Do you understand now what a negative exponent means?

OpenStudy (mathstudent55):

Use this rule to change the left side of your equation to an expression with a positive exponent.

OpenStudy (anonymous):

Is it 3?

OpenStudy (mathstudent55):

Here is your equation again. We need to follow the rule above and change just the left side to an expression with a positive exponent. Then we copy the right side just as it is. \(\large k^{-3} = \dfrac{1}{27}\) \( \large \dfrac{1}{k^3} = \dfrac{1}{27} \)

OpenStudy (anonymous):

Isn't it 3?

OpenStudy (mathstudent55):

You mean the final answer?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\(\large \dfrac{1}{k^3} = \dfrac{1}{3^3} \) \(k = 3\) You are correct.

OpenStudy (anonymous):

Thank You!

OpenStudy (mathstudent55):

You're welcome.